An ornament weighing 36 g in air, weighs only 34 g in water. Assuming that some copper is mixed with gold to prepare the ornament, find the amount of copper in it. Specific gravity of gold is 19.3 and that of copper is 8.9.

Grade:11

1 Answers

Jitender Pal
8 years ago
Sol. Here, m0= mAu+ mcu= 36 g …(1) Let V be the volume of the ornament in cm3 So, V × ρw× g = 2 × g ⇒(Vau+ Vcu) × ρw× g = 2 × g ⇒ (m/pau + m/pau) ρw × g = 2 × g ⇒ (pAu/19.3 + pAu/8.9) × 1 = 2 ⇒8.9 mAu+ 19.3 mcu= 2 × 19.3 × 8.9 = 343.54 …(2) From equation (1) and (2), 8.9 mAu+ 19.3 mcu= 343.54 ⇒ 8.9(mAu + mCu/mCu = 2.225g) So, the amount of copper in the ornament is 2.2 g.